Please help, don't understand...thanks
A small publishing company is planning to publish a new book. The production costs will include one-time fixed costs (such as editing) and variable costs (such as printing). The one-time fixed costs will amount to $35,630. The variable costs will be $11.50 per book. The publisher will sell the finished product to bookstores at a price of $20.25 per book. How many books must the publisher print and sell so that the production costs will equal the money obtained from sales?

Answer by scott8148(5895)

let n="number of books to break even"

35630+n(11.50)=n(20.25) __ 35630=n(8.75) __ n=4072